OpenStudy (anonymous):

A stone is thrown upward with a speed of 30 feet per second from the edge of a cliff 200 feet above the ground. What is the speed of the stone when it hits the ground? Use acceleration due to gravity as –32 feet per second squared and approximate your answer to 3 decimal places.

OpenStudy (anonymous):

I figured out my velocity equation which is v(t)= -32t+30 and I know that the object reaches a velocity of zero at t=30/32

OpenStudy (anonymous):

If you need me to check it yes your correct exactly on that. Time wise, it would take 0.15 seconds to get to the bottom. (-32) with in place, 0.15 seconds till rock hits bottom. Hope this helps. If you have any more questions you need help on tag me or msg me them! Have a nice day. Also cute profile picture :)

OpenStudy (anonymous):

How do you know it hits the ground at .15 seconds?

OpenStudy (anonymous):

very simple. 30/200. :)

OpenStudy (anonymous):

Since velocity equals zero at t=30/32 or .9375 seconds, the stone has only began it's descent at nearly 1 second.

OpenStudy (anonymous):

Ahh I'm so dang confused right now. Jesus though your adorable I need to say that right now lol. Anyways true zero at t=30/32 / .9375 seconds. So that would equal a range of around, what, 30 feet = 32 seconds? I'm just so confused and my whole paper is full of stupid velocity multiplication and testing the velocity rate, etc. It's about to turn into you helping me with velocity instead of me helping you out on this lol.

OpenStudy (anonymous):

adorable lol google "cute selfie" and you will find it i used it for a while myself

OpenStudy (sdfgsdfgs):

adorable or not ;) here are the kinematics eqns that can be used to solve this prob. http://zonalandeducation.com/mstm/physics/mechanics/kinematics/EquationsForAcceleratedMotion/EquationsForAcceleratedMotion.htm

OpenStudy (anonymous):

Lol i am studying this now... i have no idea myself... sorry girl lol

OpenStudy (anonymous):

Haha thank you

OpenStudy (anonymous):

np:)

OpenStudy (anonymous):

lol... i did nothing but YOURRRR WELCOMEEEEEEEE

OpenStudy (anonymous):

you're